A particular investment offers the following cash flows: none for 3 years, then $1500 per year...

  1. A particular investment offers the following cash flows: none for 3 years, then $1500 per year for 5 years, then $2000 per year for 10 years. The required return is 6%.
    1. What is the value of the investment today?
    2. What is the value of the investment in 5 years?
    3. What is the value of the investment in 8 years?
Answer #1

What is the value of the investment today

=(1500*((1-(1+6%)^(-5))/6%))/(1+6%)^3+(2000*((1-(1+6%)^(-10))/6%))/(1+6%)^8

=14540.79

b.

What is the value of the investment in 5 years?

=1500/(1+6%)^1+1500/(1+6%)^2+1500/(1+6%)^3+(2000*((1-(1+6%)^(-10))/6%))/(1+6%)^3

=16368.86

c.

What is the value of the investment in 8 years

=(2000*((1-(1+6%)^(-10))/6%))

=14720.17

the above is answer..